OpenStudy (anonymous):

use elimination to solve the system 3x-5y=8 and 4x-7y=10. what is the x value in the solution? A. -1 B. 2 C.5 D.6

OpenStudy (uri):

3x-5y=8 (muliply by -7) 4x-7y=10(muliply by 5) -21x+35y=-56 20x-35y=50 ========== -1x=-6 x=-6/-1 x=6 ans

OpenStudy (anonymous):

that was great, thank you (: